# Chaos and Synchronization in Financial Leverages Dynamics: Modeling Systemic Risk with Coupled Unimodal Maps

### v1

@@ Abstract
Financial systemic risk refers to the event in which several institutions fail simultaneously, possibly leading to large financial crisis. To model the evolution of the level of risk associated with a system, it turns out to be crucial to look at the evolution of banks’ leverage ratios, quantities that can amplify both gains and losses. Building on the works lillo2023unimodalmazzarisi2014dynamicalcorsi2016micro, we propose a simple analytical dynamical model for the evolution of coupled leverages and explore how this can lead to complex behavior within the banking system. In particular, we investigate how the heterogeneity in bank size and strategies of single banks may impact the stability of the whole system. These insights contribute to a deeper understanding of the mechanisms that can precipitate systemic crises and inform strategies for enhancing financial resilience.
@@ I Introduction
Several mechanisms may be responsible for the emergence of systemic risk in a system of financial institutions (or banks for simplicity), such as informational contagion leading to bank runs and liquidity shortages, direct contagion in which banks fail to pay back loans to other banks, etc. In a broad sense, systemic risk may also be considered to include the risk associated with a system collapsing simultaneously as a result of a shock, either endogenous or exogenous (de2000systemic). For instance, in the case in which multiple banks hold positions in the same asset, a severe loss of its value would cause all banks to fail together.
@@
The leverage $\lambda$ of a market agent, such as a bank, is defined as the ratio between the value of the assets held $A$, i.e. the economic resources the bank owns, and the value of its equity $E$, i.e. what remains of the asset value after subtracting the liabilities to which the bank is subject to. As the name itself suggests, a large leverage means an amplification of resources invested, which is made possible by the bank borrowing money to invest (a “pumping” of money from the creditors allows the bank to amplify its invested resources) and which translates into an amplification of gains, but also losses.
The relationship between leverage and risk has indeed already been suggested by many (acharya2016banks; adrian2010liquidity; adrian2014procyclical; fostel2008leverage; huizinga2012bank; nuno2017bank), and depends on the following mechanism. In order to maximize gain, banks want to be as leveraged as possible, as a larger leverage will in general allow for more opportunities to make profit by buying on margin. On the other hand, banks have to face limitations on the maximum attainable leverage imposed by financial regulations in order to make the financial system more robust and resilient.
@@
$\int_{-\infty}^{-\Lambda_{\text{VaR}}}f(r^{p})dr^{p}=P_{\text{VaR}}$
$r^{p}$ being the portfolio return and $f$ its probability density function. In other words, banks are required to loose more than $\Lambda_{\text{VaR}}$ with a probability smaller than $P_{\text{VaR}}$. Having defined $\Lambda_{\text{VaR}}$, it is required that the maximum loss, $\Lambda_{\text{VaR}}$, be smaller than the equity of the bank. This translates in a maximum leverage that banks are allowed to set, i.e. $\lambda\leq\frac{1}{\Lambda_{\text{VaR}}}$.
Notice that $\Lambda_{\text{VaR}}$ depends on the variance of $r^{p}$, i.e. on the volatility of asset prices. For example, for Gaussian returns with variance $\sigma^{2}$, a probability of loss $P_{\text{VaR}}=5\%$ would give $\Lambda_{\text{VaR}}\approx 1.64\sigma$. Banks will thus make an estimate of this variance and set a target leverage $\lambda=\frac{1}{1.64\sigma}$.
This leverage-targeting strategy banks may adopt can result, due to the finite liquidity of the market, in a positive feedback between leverages and asset prices (adrian2010liquidity), possibly leading to instabilities very much like what happens in physical systems with positive feedback. Being concrete, what happens is that if things go well, in order to maintain a certain leverage level, banks are led to borrow more; while if things go bad banks will start selling assets. For example, consider a bank having a target leverage $\lambda=10$, which it realizes having an asset value equal to $100$ and an equity equal to $10$. If there is a $1\%$ portfolio return, at the next time the bank will have an asset value of $101$ and an equity equal to $11$, with a resulting leverage $\frac{101}{11}\approx 9.2$, which is less than the target leverage. Thus, to increase its leverage, the bank will borrow some money and increase its asset value, determining as a consequence an increase in its price.
This is also known as procyclicality of the Value-at-Risk constraint and has the effect of amplifying prices movements, which is particularly relevant during a falling period of a financial crisis.
@@
In the case in which the banks are “similar” in the sense that they have comparable sizes and risk estimation strategies, one can reduce to study the evolution of a “representative” leverage. This is what has been done for instance in a paper by Lillo et al.lillo2023unimodal, where they obtain a slow-fast deterministic-random dynamical system for the leverage of a bank investing in a single asset. We take this model as a starting point, and introduce heterogeneity in the system, allowing the banks to have different asset sizes and different risk estimation strategies. In this case, disregarding random fluctuations, the system can be modeled with a discrete time dynamical system consisting of unimodal maps coupled through a mean field (see equation 1 below).
In most previous works considering the impact of risk management practices on the dynamical properties of the leverages, either a single bank was considered (lillo2023analysis; aymanns2016taming) or multiple banks and multiple assets (corsi2016micro; mazzarisi2019panic) but with all banks being in some sense equivalent (they all solved the same optimization problem and their portfolios were determined by random choices in a pool of equivalent assets; thus, in particular, their leverages were at all instants equal). The models introduced here try to go towards a more realistic direction by differentiating banks on the basis of their capital size and their risk forecasting strategies.
Other works related to the dynamical modeling of systemic risk, following some different approaches, can be found in poledna2014leverage; castellacci2015modeling; choi2012financial; choi2013financial; geanakoplos2010leverage; awiszus2022market; capponi2020swing; thurner2011systemic. However, for example, poledna2014leverage and thurner2011systemic consider an agent-based model in which the agents don’t determine their choices on the basis of a strategy (e.g. by solving an optimization problem) and are endowed with infinite capital, while capponi2020swing and awiszus2022market don’t take into account the feedback mechanism due to the estimations of risk by the banks.
Organization of the paper. First, the case in which there is only one bank investing in a single asset (that is the case studied in lillo2023analysis; lillo2023unimodal) is recalled. Then, the focus will be on the case in which there are two banks investing in a single common asset. Both numerical simulations and analytical results (for some special configurations) are provided.
@@ II Unimodal Evolution of the Leverage for One Bank Trading One Asset
In the case in which there is only one bank investing in a single asset, the evolution of its leverage $\lambda_{t}$, with $t\in\mathbb{N}$ (in the limit in which the time scale for the financial transactions is much faster than the one used to adjust the target leverage) will turn out to be describedlillo2023unimodal by the map $T:[0,1+\gamma]\rightarrow\mathbb{R}$:
$T(x)=\left(\frac{\omega}{x^{2}}+\frac{(1-\omega)\alpha^{2}\gamma^{2}\Sigma_{\epsilon}}{(1+\gamma-x)^{2}}\right)^{-\frac{1}{2}}$
@@
Indeed, it is interesting to study how the behavior changes as $\omega$ is varied, since this is a parameter the bank can directly control. As for the values of the other parameters employed, we follow mazzarisi2014dynamical by setting $\alpha=1.64$ and $\gamma=100$; $\Sigma_{\epsilon}$ has been put equal to $0.0015^{2}$ (thus smaller than the one in the cited reference, as it is nevertheless to be expected since there only the map obtained in the $\gamma\gg 1$ is simulated, which is different from $T$). The choice of this value is not crucial, as the behavior of the system does not qualitatively change for small variations of it. However, for significantly bigger values of this parameter (e.g. $\gtrsim 0.015^{2}$) the behavior becomes trivial and therefore not so interesting to investigate, with the leverage reaching a fixed point for every value of $\omega\in[0,1]$. It is also the case that, for these choices of parameters (and for $\omega$ in a subinterval of $[0,1]$), one can find an invariant interval in $1,1+\gamma]$ and can thus guarantee that $\lambda_{t}\in[1,1+\gamma]$ for every $t\in\mathbb{N}$, as required by the interpretation of the quantity $\lambda_{t}$ as a financial leverage (see the supplementary material for more details on this). Lastly, these choices will be the “standard ones” in numerical simulations in the following sections as well. Coming back to the simulations, it is observed that for smaller values of the memory $\omega$ the system becomes more unstable, as can also be seen by looking at the Lyapunov of the map $T$ varying $\omega$ (see Fig. 3)
@@ III Coupled Unimodal Maps Model for the Leverage Evolution of Banks Trading a Common Asset
$\lambda_{i,t}=\left(\frac{\omega_{i}}{\lambda_{i,t-1}^{2}}+\frac{(1-\omega_{i})\alpha^{2}\gamma^{2}\Sigma_{\epsilon}}{\left(1+\gamma-\sum_{i=1}^{2}\pi_{i}\lambda_{i,t-1}\right)^{2}}\right)^{-\frac{1}{2}}.$ (1)
@@
For $N=2$ and in the “homogeneous case” in which $\omega_{1}=\omega_{2}$, the two leverages will completely synchronize (see Section III.2), and this happens regardless of $\pi_{1}$.
@@
Considering now the case in which $\pi_{1}\neq 0,1$, interesting bifurcation diagrams for the leverages of the two banks are observed as relevant parameters are changed. In particular, by keeping $\omega_{1}$ and $\omega_{2}$ fixed and changing the weight $\pi_{1}$ one observes a bifurcation diagram which “connects” the behaviors one would have in the forced-forcing systems corresponding to $\pi_{1}=0$ (where only $\omega_{2}$ has a role in determining the nature of the behavior) and $\pi_{1}=1$ (where now $\omega_{1}$ is the relevant parameter). Therefore, for example, there exist values for $\omega_{1}$, $\omega_{2}$ for which there is a threshold value for $\pi_{1}$ above which both banks switch from a regular behavior to an aperiodic one (see Fig. 5)
Figure 5: Example of asymptotic orbits for the two banks as $\pi_{1}$ varies and for different choices of $\omega_{1},\omega_{2}$. Here $\omega_{1}=0.5,\omega_{2}=0.3$. In doing the plot, the first 1000 values have been discarded and the next 500 plotted.
@@
Figure 6: Example of asymptotic orbits for the two banks as $\omega_{1}$ varies for $\pi_{1}=0.5$ and $\omega_{2}=0.4$. In doing the plot, the first 1000 values have been discarded and the next 800 plotted.
@@ III.1 The Model
In the case of a single bank, referring to the model introduced in corsi2016micro and further developed in lillo2023unimodal; mazzarisi2019panic, the evolution of its leverage will be the result of the following main principles: 1) to increase profit, a bank wants to maximize its leverage up to the limit imposed by regulations; 2) this limit depends on the risk associated to the asset that is estimated by the bank in a way that can vary from bank to bank; 3) the time evolution of the prices of the asset is characterized by an autoregressive process with a time scale faster than the one in which the banks update their leverage; 4) the relative size of the total assets detained by the banks (i.e. their weights $\pi_{i}$) are approximately constant.
@@
This results in a slow-fast deterministic-random dynamical system with heteroschedastic noise. In the limit in which the fast time scale is much faster than the slow one, this reduces to a deterministic dynamical system given in (1). We now show how to derive this expression by generalising the derivation given in lillo2023analysis to the case of multiple banks.
@@
Now consider how fixing a target leverage impacts the movement of the asset’s price. Having set a value for the target leverage, the bank will trade the asset between time $t$ and $t+1$ in order to keep its leverage equal to this target value. This trading process takes places at a faster pace than the one in which the target leverage is updated (i.e. at every integer time). Let’s thus introduce a quantity $\mathbb{n}\in\mathbb{N}$ so that trading operations occur $\mathbb{n}$ times every integer time step. Next, call $r_{t+\frac{k}{\mathbb{n}}}$, with $k=1,2,\dots,\mathbb{n}$ the returns on investments (i.e. the ratio between the gain/loss yielded by the investment and the initial cost of the investment). The dynamics of these returns can be thought of as made of two components:
$r_{t+\frac{k}{\mathbb{n}}}=\epsilon_{t+\frac{k}{\mathbb{n}}}+e_{t+\frac{k-1}{\mathbb{n}}};$
the first term in the RHS is an exogenous component (depending on external, non modeled events) given by a white noise term with variance $\sigma^{2}_{\epsilon}$, while the second term in the RHS is instead an endogenous component, which is the one depending on the trading actions of the bank. More in detail, the endogenous component $e_{t+h/\mathbb{n}}$ depends on the demand for the asset arising from portfolio rebalancing of the bank. Given a fractional time $s=t+\frac{h}{\mathbb{n}},h=1,2,\dots,\mathbb{n}$, let’s call the desired asset size for a generic bank $i$ $A^{\star}_{i,s}\doteq\lambda_{i,t}E_{i,s}$. Therefore at $s$ bank $i$ rebalances its portfolio by trading the quantity
$A^{\star}_{i,s}-A_{i,s}$ $=\lambda_{i,t}E_{i,s}-A^{\star}_{i,s-\frac{1}{\mathbb{n}}}(1+r_{s})$
$=\lambda_{i,t}(E_{i,s-\frac{1}{\mathbb{n}}}+r_{s}A^{\star}_{i,s-\frac{1}{\mathbb{n}}})-A^{\star}_{i,s-\frac{1}{\mathbb{n}}}(1+r_{s})$
$=(\lambda_{i,t}-1)r_{s}A^{\star}_{i,s-\frac{1}{\mathbb{n}}}$
The total demand for the asset will be given by
$D_{s}=\sum_{i=1}^{N}(A^{\star}_{i,s}-A_{i,s})$
Assuming a standard linear price impact function, the endogenous component $e_{s}$ for the return of the asset is given by
$e_{s}=\frac{D_{s}}{\gamma C_{s}}$
@@
$C_{s}=\sum_{i=1}^{N}A^{\star}_{i,s-\frac{1}{\mathbb{n}}}$
is a proxy for market capitalization of the investment. Thus:
$r_{s}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}-1)A^{\star}_{i,s-\frac{2}{\mathbb{n}}}}{\gamma\sum_{i=1}^{N}A^{\star}_{i,s-\frac{2}{\mathbb{n}}}}r_{s-\frac{1}{\mathbb{n}}}+\epsilon_{s}$
@@
Now notice that (recall $A^{\star}_{i,s}\doteq\lambda_{i,t}E_{i,s}$ and $E_{i,s}=\prod_{q=0}^{s}(1+r_{q})E_{i,s}$) if the leverages were the same for all banks, then $\pi_{i,s}$ would not depend on the time $s$. Thus, in the homogeneous case we can consider the weights as independent of $s$. Moreover, numerical simulations show that the error made by considering the weights as independent of $s$ even in the heterogeneous case is small enough to make it reasonable to start by analyzing the simplest case in which the weights are constant in time. All the more reason for the fact that for the cases studied analytically more in depth (i.e. the homogeneous case and the forced-forcing one) this approximation is exact. Let’s therefore put $\pi_{i}=\pi_{i,s}$ and look at $\pi_{i}$ as a measure of the “size” of bank $i$. It is then possible to write:
$r_{s}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}\pi_{i}-1)}{\gamma}r_{s-\frac{1}{\mathbb{n}}}+\epsilon_{s}$

### v2

@@ Abstract
Financial systemic risk (or systemic failure) refers to the event in which several institutions fail simultaneously, possibly leading to large financial crisis. To model the evolution of the level of risk associated with a system, it turns out to be crucial to look at the evolution of banks’ leverage ratios, quantities that can amplify both gains and losses. Building on the works Lillo et al. (2023b)Mazzarisi et al. (2019)Corsi et al. (2016), we propose a simple analytical dynamical model for the evolution of coupled leverages and explore how this can lead to complex behavior within the banking system. In particular, we investigate how the heterogeneity in bank size and strategies of single banks may impact the stability of the whole system. These insights contribute to a deeper understanding of the mechanisms that can precipitate systemic crises and inform strategies for enhancing financial resilience.
@@ I Introduction
Several mechanisms may be responsible for the emergence of systemic risk in a system of financial institutions (or banks for simplicity), such as informational contagion leading to bank runs and liquidity shortages, direct contagion in which banks fail to pay back loans to other banks, etc. In a broad sense, systemic risk may also be considered to include the risk associated with a system collapsing simultaneously as a result of a shock, either endogenous or exogenous (De Bandt and Hartmann (2000)). For instance, in the case in which multiple banks hold positions in the same asset, a severe loss of its value would cause all banks to fail together.
@@
The leverage $\lambda$ of a market agent, such as a bank, is defined as the ratio between the value of the assets held $A$, i.e. the economic resources the bank owns, and the value of its equity $E$, i.e. what remains of the asset value after subtracting the liabilities $L$ to which the bank is subject to, for instance (so that E=ALE=A-L). As the name itself suggests, a large leverage means an amplification of resources invested, which is made possible by the bank borrowing money to invest (a “pumping” of money from the creditors allows the bank to amplify its invested resources) and which translates into an amplification of gains, but also losses.
The relationship between leverage and risk has indeed already been suggested by many (Acharya and Ryan (2016); Adrian and Shin (2010, 2014); Fostel and Geanakoplos (2008); Huizinga and Laeven (2012); Nuno and Thomas (2017)), and depends on the following mechanism. In order to maximize gain, banks want to be as leveraged as possible, as a larger leverage will in general allow for more opportunities to make profit by buying on margin (i.e. by investing borrowed money). On the other hand, banks have to face limitations on the maximum attainable leverage imposed by financial regulations in order to make the financial system more robust and resilient.
@@
$\int_{-\infty}^{-\Lambda_{\text{VaR}}}f(r^{p})dr^{p}=P_{\text{VaR}}$ (1)
$r^{p}$ being the portfolio return (i.e. the ratio between the gain/loss yielded by the investment and the initial cost of the investment) and $f$ its probability density function. In other words, banks are required to assure that the probability of losing an amount larger than $\Lambda_{\text{VaR}}$ is less than $P_{\text{VaR}}$. Having defined $\Lambda_{\text{VaR}}$, it is required that the maximum loss, $A\Lambda_{\text{VaR}}$, be smaller than the equity of the bank. This translates in a maximum leverage that banks are allowed to set, i.e. $\lambda\leq\frac{1}{\Lambda_{\text{VaR}}}$. Assume a mean $\mu$ and a variance $\sigma^{2}$ for $r^{p}$. Then, by Chebychev inequality we can write:
$\frac{\sigma^{2}}{(\mu+\Lambda_{\text{VaR}})^{2}}\geq P(|r^{p}-\mu|\geq\mu+\Lambda_{\text{VaR}})\geq P(r^{p}\leq-\Lambda_{\text{VaR}})$ (2)
So that
$P(r^{p}\leq-\Lambda_{\text{VaR}})\leq\frac{\sigma^{2}}{(\mu+\Lambda_{\text{VaR}})^{2}}$ (3)
Requiring $\frac{\sigma^{2}}{(\mu+\Lambda_{\text{VaR}})^{2}}\leq P_{\text{VaR}}$, i.e. $\Lambda_{\text{VaR}}\geq\frac{\sigma}{\sqrt{P_{\text{VaR}}}}-\mu$, ensures that $P(r^{p}\leq-\Lambda_{\text{VaR}})\leq P_{\text{VaR}}$. We deduce the following upper bound for the leverage: $\lambda\leq\frac{1}{\Lambda_{\text{VaR}}}\leq\frac{1}{\frac{\sigma}{\sqrt{P_{\text{VaR}}}}-\mu}$ Actually, the most stringent inequality is usually required to hold: $\lambda\leq\frac{1}{\Lambda_{\text{VaR}}}\leq\frac{\sqrt{P_{\text{VaR}}}}{\sigma}$
Banks will thus make an estimate of $\sigma$ and set a target leverage $\lambda=\frac{1}{\alpha\sigma}$. Having defined $\alpha=\frac{1}{\sqrt{P_{\text{VaR}}}}$. This leverage-targeting strategy banks may adopt can result in a positive feedback between leverages and asset prices (Adrian and Shin (2010)), possibly leading to instabilities very much like what happens in physical systems with positive feedback. Being concrete, what happens is that if things go well (i.e. the price of the asset rises and thus $E$ increases), in order to maintain a certain leverage level, banks must increase $A$, and are thus led to borrow more (i.e. increase $L$); while if things go bad (price falls and thus $E$ decreases) banks will have to reduce $A$ (by reducing $L$). Now, the price of the asset increases (decreases) as the demand for it increases (decreases). Thus if price rise (fall), banks will increase (reduce) $A$ to reach the target leverage, i.e. the demand for the asset will rise (fall) and so will do its price, making apparent the positive feedback between asset price and leverages.
For example, consider a bank having a target leverage $\lambda=10$, which it realizes having an asset value equal to $100$ and an equity equal to $10$. If there is a $1\%$ portfolio return, at the next time the bank will have an asset value of $101$ and an equity equal to $11$, with a resulting leverage $\frac{101}{11}\approx 9.2$, which is less than the target leverage. Thus, to increase its leverage, the bank will borrow some money and increase its asset value, determining as a consequence an increase in its price.
This positive feedback mechanism is also known as procyclicality of the Value-at-Risk constraint Adrian and Shin (2014) and has the effect of amplifying prices movements, which is particularly relevant during a falling period of a financial crisis (i.e. when prices of assets drop).
@@
Indeed, the larger the leverages, the more amplified asset-price movements are (due to the first feedback mechanism mentioned above) and thus the riskier it is for the price to dramatically fall (maybe after a negative fluctuation), making all banks investing in it to fail simultaneously, i.e. to an increase in the level of systemic risk.
In the case in which the banks are “similar” in the sense that they have comparable sizes and risk estimation strategies, one can reduce to study the evolution of a “representative” leverage. This is what has been done for instance in a paper by Lillo et al.Lillo et al. (2023b), where they obtain a slow-fast deterministic-random dynamical system for the leverage of a bank investing in a single asset. We take this model as a starting point, and introduce heterogeneity in the system, allowing the banks to have different asset sizes and different risk estimation strategies. In this case, disregarding random fluctuations, the system can be modeled with a discrete time dynamical system consisting of unimodal maps coupled through a mean field (see equation 5 below).
In most previous works considering the impact of risk management practices on the dynamical properties of the leverages, either a single bank was considered (Lillo et al. (2023a); Aymanns et al. (2016)) or multiple banks and multiple assets (Corsi et al. (2016); Mazzarisi et al. (2019)) but with all banks being in some sense equivalent (they all solved the same optimization problem and their portfolios were determined by random choices in a pool of equivalent assets; thus, in particular, their leverages were at all instants equal). The models introduced here try to go towards a more realistic direction by differentiating banks on the basis of their capital size and their risk forecasting strategies.
Other works related to the dynamical modeling of systemic risk, following some different approaches, can be found in Poledna et al. (2014); Castellacci and Choi (2015); Choi and Douady (2012); Choi et al. (2013); Geanakoplos (2010); Awiszus et al. (2022); Capponi et al. (2020); Thurner (2011). However, for example, Poledna et al. (2014) and Thurner (2011) consider an agent-based model in which the agents don’t determine their choices on the basis of a strategy (e.g. by solving an optimization problem) and are endowed with infinite capital, while Capponi et al. (2020) and Awiszus et al. (2022) don’t take into account the feedback mechanism due to the estimations of risk by the banks.
Organization of the paper. First, the case in which there is only one bank investing in a single asset (that is the case studied in Lillo et al. (2023a, b)) is recalled. Then, the focus will be on the case in which there are two banks investing in a single common asset. Both numerical simulations and analytical results (for some special configurations) are provided.
@@ II Unimodal Evolution of the Leverage for One Bank Trading One Asset
In the case in which there is only one bank investing in a single asset, the evolution of its leverage $\lambda_{t}$, with $t\in\mathbb{N}$ (in the limit in which the time scale for the financial transactions is much faster than the one used to adjust the target leverage) will turn out to be describedLillo et al. (2023b) by the map $T:[0,1+\gamma]\rightarrow\mathbb{R}$:
$T(x)=\left(\frac{\omega}{x^{2}}+\frac{(1-\omega)\alpha^{2}\gamma^{2}\Sigma_{\epsilon}}{(1+\gamma-x)^{2}}\right)^{-\frac{1}{2}}$ (4)
@@
Indeed, it is interesting to study how the behavior changes as $\omega$ is varied, since this is a parameter the bank can directly control. As for the values of the other parameters employed, we follow Mazzarisi (2014) by setting $\alpha=1.64$ and $\gamma=100$; $\Sigma_{\epsilon}$ has been put equal to $0.0015^{2}$ (thus smaller than the one in the cited reference, as it is nevertheless to be expected since there only the map obtained in the $\gamma\gg 1$ is simulated, which is different from $T$). The choice of this value is not crucial, as the behavior of the system does not qualitatively change for small variations of it. However, for significantly bigger values of this parameter (e.g. $\gtrsim 0.015^{2}$) the behavior becomes trivial and therefore not so interesting to investigate, with the leverage reaching a fixed point for every value of $\omega\in[0,1]$. It is also the case that, for these choices of parameters (and for $\omega$ in a subinterval of $[0,1]$), one can find an invariant interval in $1,1+\gamma]$ and can thus guarantee that $\lambda_{t}\in[1,1+\gamma]$ for every $t\in\mathbb{N}$, as required by the interpretation of the quantity $\lambda_{t}$ as a financial leverage (see the supplementary material for more details on this). Lastly, these choices will be the “standard ones” in numerical simulations in the following sections as well. Coming back to the simulations, it is observed that for smaller values of the memory $\omega$ the system becomes more unstable, as can also be seen by looking at the Lyapunov of the map $T$ varying $\omega$ (see Fig. 3)
@@ III Coupled Unimodal Maps Model for the Leverage Evolution of Banks Trading a Common Asset
$\lambda_{i,t}=\left(\frac{\omega_{i}}{\lambda_{i,t-1}^{2}}+\frac{(1-\omega_{i})\alpha^{2}\gamma^{2}\Sigma_{\epsilon}}{\left(1+\gamma-\sum_{i=1}^{2}\pi_{i}\lambda_{i,t-1}\right)^{2}}\right)^{-\frac{1}{2}}$ (5)
@@
Remark: We are interested in considering the behavior of the system for different values of the weights of the banks in terms of asset size $\pi_{1}$ and $\pi_{2}$, and of the “memories” of the banks $\omega_{1}$, and $\omega_{2}$. In other words, for our purposes the bifurcation parameters are $\omega_{1},\omega_{2},\pi_{1}$, while all others parameters are intended fixed.
For $N=2$ and in the “homogeneous case” in which $\omega_{1}=\omega_{2}$, the two leverages will completely synchronize, as in Fig. 4 (see Section III.2), and this happens regardless of $\pi_{1}$.
@@
Remark: We say that two banks are synchronised if the values of their leverages coincide. In this sense, a system of banks is synchronized if it sits on the “diagonal" of the phase space. In the situations where we say that the leverages synchronise, we mean that the diagonal is locally stable. An initial condition in the proximity of the diagonal will synchronize if it will approach asymptotically the diagonal or, equivalently, if the orbits of different banks asymptotically converge. Given this characterisation of synchronisation, a measurament of synchronisation could be the distance from the diagonal of the space.
@@
Considering now the case in which $\pi_{1}\neq 0,1$, interesting bifurcation diagrams for the leverages of the two banks are observed as relevant parameters are changed. In particular, by keeping $\omega_{1}$ and $\omega_{2}$ fixed and changing the weight $\pi_{1}$ one observes a bifurcation diagram which “connects” the behaviors one would have in the forced-forcing systems corresponding to $\pi_{1}=0$ (where only $\omega_{2}$ has a role in determining the nature of the behavior) and $\pi_{1}=1$ (where now $\omega_{1}$ is the relevant parameter). Therefore, for example, there exist values for $\omega_{1}$, $\omega_{2}$ for which there is a threshold value for $\pi_{1}$ below which both banks switch from a regular behavior to an aperiodic one (see Fig. 5)
Figure 5: Example of asymptotic orbits for the two banks (green for Bank 1 and blue for Bank 2) as $\pi_{1}$ varies and for different choices of $\omega_{1},\omega_{2}$. Here $\omega_{1}=0.8,\omega_{2}=0.3$. In doing the plot, the first 1000 values have been discarded and the next 500 plotted.
@@
Figure 6: Example of asymptotic orbits for the two banks (green for Bank 1 and blue for Bank 2) as $\omega_{1}$ varies for $\pi_{1}=0.5$ and $\omega_{2}=0.6$. In doing the plot, the first 1000 values have been discarded and the next 800 plotted.
@@ III.1 The Model
In the case of a single bank, referring to the model introduced in Corsi et al. (2016) and further developed in Lillo et al. (2023b); Mazzarisi et al. (2019), the evolution of its leverage will be the result of the following main principles: 1) to increase profit, a bank wants to maximize its leverage up to the limit imposed by regulations; 2) this limit depends on the risk associated to the asset that is estimated by the bank in a way that can vary from bank to bank; 3) the time evolution of the prices of the asset is characterized by an autoregressive process with a time scale faster than the one in which the banks update their leverage; 4) the relative size of the total assets detained by the banks (i.e. their weights $\pi_{i}$) are approximately constant.
@@
[…]
@@
Now consider how fixing a target leverage impacts the movement of the asset’s price. Having set a value for the target leverage, the bank will trade the asset between time $t$ and $t+1$ in order to keep its leverage equal to this target value. This trading process takes places at a faster pace than the one in which the target leverage is updated (i.e. at every integer time). Let’s thus introduce a quantity $\mathbb{n}\in\mathbb{N}$ so that trading operations occur $\mathbb{n}$ times every integer time step. Next, call $r_{t+\frac{k}{\mathbb{n}}}$, with $k=1,2,\dots,\mathbb{n}$ the returns on investments (as already defined in the introduction). The dynamics of these returns can be thought of as made of two components:
$r_{t+\frac{k}{\mathbb{n}}}=\epsilon_{t+\frac{k}{\mathbb{n}}}+e_{t+\frac{k-1}{\mathbb{n}}}$ (7)
the first term in the RHS is an exogenous component (depending on external, non modeled events) given by a white noise term with variance $\sigma^{2}_{\epsilon}$, while the second term in the RHS is instead an endogenous component, which is the one depending on the trading actions of the bank. More in detail, the endogenous component $e_{t+h/\mathbb{n}}$ depends on the demand for the asset arising from portfolio rebalancing of the bank, as will now be explained. Given a fractional time $s=t+\frac{h}{\mathbb{n}},h=1,2,\dots,\mathbb{n}$, let’s call the desired asset size for a generic bank $i$ $A^{\star}_{i,s}\doteq\lambda_{i,t}E_{i,s}$. Therefore at $s$ bank $i$ rebalances its portfolio by trading (i.e. buying or selling, depending on the sign of $A^{\star}_{i,s}-A_{i,s}$) the quantity
{align}
A^⋆_i,s-A_i,s=λ_i,tE_i,s-A^⋆_i,s-1n(1+r_s)
=λ_i,t(E_i,s-1n+r_s A^⋆_i,s-1n)-A^⋆_i,s-1n(1+r_s)
=(λ_i,t-1)r_s A^⋆_i,s-1n
Where it is assumed that from $t$ to $t+1$ the liabilities of the banks are constant, so that the change in their total assets derives only from that of the equity. The total demand for the asset will then be given by
$D_{s}=\sum_{i=1}^{N}(A^{\star}_{i,s}-A_{i,s})$ (8)
Assuming a standard linear price impact function (i.e. assuming that the change in the price of the asset is directly proportional to the demand for it), the endogenous component $e_{s}$ for the return of the asset is given by
$e_{s}=\frac{D_{s}}{\gamma C_{s}}$ (9)
@@
$C_{s}=\sum_{i=1}^{N}A^{\star}_{i,s-\frac{1}{\mathbb{n}}}$ (10)
is a proxy for market capitalization of the investment. What we are saying here is that the price change is larger when the ratio of the demand for it to the total amount of money invested in the asset is large and when the liquidity is low (indeed, by definition the liquidity measures the impact of selling/buying an asset on its price).
Thus:
$r_{s}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}-1)A^{\star}_{i,s-\frac{2}{\mathbb{n}}}}{\gamma\sum_{i=1}^{N}A^{\star}_{i,s-\frac{2}{\mathbb{n}}}}r_{s-\frac{1}{\mathbb{n}}}+\epsilon_{s}$ (11)
@@
Now notice that (recall $A^{\star}_{i,s}\doteq\lambda_{i,t}E_{i,s}$ and $E_{i,s}=\prod_{q=0}^{s}(1+r_{q})E_{i,s}$) if the leverages were the same for all banks, then $\pi_{i,s}$ would not depend on time; this is a consequence of the fact that all banks are investing in the same asset. More precisely, the weights in this case would be constant both within a unit time interval (e.g. between t1t-1 and $t$) and across different times $t$. Thus, in the homogeneous case we can consider the weights as independent of $s$. Moreover, numerical simulations show that the error made by considering the weights as independent of $s$ even in the heterogeneous case is small enough to make it reasonable to start by analyzing the simplest case in which the weights are constant in time, leaving the investigation of the general case in which the weights change over time to future works. All the more reason for the fact that for the cases studied analytically more in depth (i.e. the homogeneous case and the forced-forcing one) this approximation is exact. Let’s therefore put $\pi_{i,s}=\pi_{i}$ for every $s\in\frac{\mathbb{N}}{\mathbb{n}}$ and look at $\pi_{i}$ as a measure of the “size” of bank $i$. It is then possible to write:
$r_{s}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}\pi_{i}-1)}{\gamma}r_{s-\frac{1}{\mathbb{n}}}+\epsilon_{s}$ (12)